OpenStudy (blackstreet23):

Sketch the curve in polar coordinates r^2 = 16 sin 2(theta)

OpenStudy (blackstreet23):

@Hero @Shalante @IrishBoy123

hero (hero):

It's best to use a polar graph to plot the points

OpenStudy (blackstreet23):

umm I think i might know what i have to do

OpenStudy (blackstreet23):

but i just dont know is that is right

OpenStudy (blackstreet23):

I mean i trying to find the endpoints to just draw a rought sketch

OpenStudy (blackstreet23):

I got that \[r = +or- (r*\sqrt{2\sin(\theta)}\]

OpenStudy (blackstreet23):

sorry 4

OpenStudy (blackstreet23):

4 * root(sin2(theta))

OpenStudy (blackstreet23):

?

hero (hero):

The graph is the same for both positive and negative expressions

OpenStudy (blackstreet23):

ohh ok. I think i get it